Illustrate the graph of the given one-to-one function and its inverse.
Activity 13.2
1. f(x) = 3x - 4
2. f(x) = 5x + 3
3. f(x) = 7-5
4. f(x)
x-2
3
5. f(x)
x + 3
2​
1
Expert's answer
2021-10-08T08:01:41-0400

1 "f(x)=3x-4 \implies f^{-1}(x)=\frac{x+4}{3}"

Below is the graph of "f(x)"(color green) and "f^{-1}(x)" (color red)



2 "f(x)=5x+3 \implies f^{-1}(x)=\frac{x-3}{5}"

Below is the graph of "f(x)"(color green) and "f^{-1}(x)" (color red)



3 "f(x)=7x-5 \implies f^{-1}(x)=\frac{x+5}{7}"

Below is the graph of "f(x)"(color green) and "f^{-1}(x)" (color red)



4 "f(x)=\frac{x-2}{3} \implies f^{-1}(x)=3x+2"

Below is the graph of "f(x)"(color green) and "f^{-1}(x)" (color red)



5 "f(x)=\frac{x+3}{2} \implies f^{-1}(x)=2x-3"

Below is the graph of "f(x)"(color green) and "f^{-1}(x)" (color red)